# Explainable AI for Comparative Analysis of Intrusion Detection Models

### v1

@@ I Introduction
  • Random Forest shows the best robust by treating all input features equally, and therefore, it still maintains the best performance while top features are not available.
@@ III-A Data Preprocessing
Figure 2: Feature correlation matrix
Figure 3: Selected features for binary classifiers.
Figure 4: Selected features for multi-class classifiers.
Figure 5: Feature sensitivity of binary intrusion detection model classifiers trained with complete features.
@@ IV-A Binary Classifiers
The feature sensitivity with respect to classification accuracy degradation of binary classifiers are given in Figures 5 and  6. As depicted, most binary models are extremely sensitive to less than three top features, in particular, Decision Tree model is extremely sensitive to even single feature occulusion occlusion. Meanwhile, Multi-Layer Feed-Forward Neural Network and Random Forest models exploit more features than other models. We adjusted the L2 regularization coefficient of MLP model from 0.0001 to 0 and we did not observe significant differences. A possible explanation is that the neural network only leverages a few highly important features and thus develops a sparse internal structure which is not sensitive to the L2 regularization.
@@ IV-B Multi-Class Classifiers
Figure 10: Feature sensitivity of binary intrusion detection model classifiers trained with complete features.

### v2

@@ I Introduction
  • Random Forest show the best robust by treating all input features equally, and therefore, it still maintains the best performance while top features are not available.
@@ II Related Work
Figure 2: Feature correlation matrix
Figure 3: Selected features for binary classifiers.
Figure 4: Selected features for multi-class classifiers.
Figure 5: Feature sensitivity of intrusion detection model classifiers trained with complete features.
@@ IV-A Binary Classifiers
The feature sensitivity with respect to classification accuracy degradation of binary classifiers are given in Figures 5 and  6. As depicted, most binary models are extremely sensitive to less than three top features, in particular, Decision Tree model is extremely sensitive to even single feature occlusion. Meanwhile, Multi-Layer Feed-Forward Neural Network and Random Forest models exploit more features than other models. We adjusted the L2 regularization coefficient of MLP model from 0.0001 to 0 and we did not observe significant differences. A possible explanation is that the neural network only leverages a few highly important features and thus develops a sparse internal structure which is not sensitive to the L2 regularization.
@@ IV-B Multi-Class Classifiers
Figure 9: Feature sensitivity of intrusion detection model classifiers trained without the top features.